OpenStudy (anonymous):

What is the value of x? 6(-3-x)-2x=14

OpenStudy (anonymous):

@bibby

pooja195 (pooja195):

Start by distributing \[\rm~(6)(−3)+(6)(−x)−2x=14\]

OpenStudy (anonymous):

ok

pooja195 (pooja195):

6 x -3= ?

OpenStudy (anonymous):

-18

pooja195 (pooja195):

6 x -X =?

OpenStudy (anonymous):

im confused

pooja195 (pooja195):

On what?

OpenStudy (anonymous):

the last reply you put up

pooja195 (pooja195):

6 times -x put the 2 together to get -6x

pooja195 (pooja195):

\[\huge~\rm~−18−6x−2x=14\] Combine like terms \[\huge~\rm(−6x−2x)+(−18)=14\]

OpenStudy (hitaro9):

Maybe instead of using x for multiplication, we could stick to using the parenthesis. Have you seen this notation before? 6(3) = 18?

pooja195 (pooja195):

-6-2=?

OpenStudy (anonymous):

-8

OpenStudy (anonymous):

and yes @Hitaro9

pooja195 (pooja195):

\[\huge~\rm~−8x−18=14\] Now add 18 to both sides. 14+18=?

OpenStudy (anonymous):

32

pooja195 (pooja195):

good -8x=32 divide both sides by -8 :) 32/-8=?

OpenStudy (anonymous):

-4

pooja195 (pooja195):

Good so you get x=-4 :)

OpenStudy (hitaro9):

So were you having trouble with how to solve 6(-3-x) ?

OpenStudy (anonymous):

Thank you @pooja195

pooja195 (pooja195):

np :) @Hitaro9 I think they understand. It was my mistake I shouldn't have put the x's as multiplication signs.

OpenStudy (anonymous):

No @Hitaro9, I was just confused with poojas response

hartnn (hartnn):

factoring and distributing \(\large ☻(☺+♥) = ☻☺+☻♥ \\ \text {this is distributing}\) \(\large ☻☺+☻♥= ☻(☺+♥) \\ \text{this is factoring, } \\ \text{☻ was common from both terms and factored out.}\)

OpenStudy (hitaro9):

Alright. What was snagging you originally? Why'd you need to come to openstudy for help?

OpenStudy (anonymous):

trying to figure out what the value of x was in a bunch of problems.

pooja195 (pooja195):

If you have more feel free to tag me i'll help you out! :)

hartnn (hartnn):

mataji, it'd be better if you show your attempt first :) Thanks.

OpenStudy (anonymous):

Ok, Thanks again @pooja195

pooja195 (pooja195):

yw

OpenStudy (hitaro9):

Right, that was the goal of the problem, what was keeping you from getting to it?

OpenStudy (anonymous):

Trying to figure out how to do the main problem

OpenStudy (hitaro9):

So are you uncertain about variables in general? If I asked you to solve x+3 =5, would you be able to solve for x?

OpenStudy (anonymous):

What is the value of x? 10x+4=12x-16

OpenStudy (anonymous):

Yes. Very confused in general

pooja195 (pooja195):

Have you attempted it? Can you show your steps ? Show what you know :)

hartnn (hartnn):

heard of "combining like terms" ? first find out which terms are "like"

OpenStudy (hitaro9):

Okay. Rather than get bogged down in these larger problems, let's focus on an easier problem so we can get the concepts

OpenStudy (anonymous):

No @Hitaro9

OpenStudy (anonymous):

That was to your last question sorry

OpenStudy (anonymous):

What is the value of x? 10x+4=12x-16

OpenStudy (hitaro9):

Okay so Durgamata. Can you tell me about what a variable is? What can you tell me about x?

OpenStudy (anonymous):

Umm, X is a letter. LOL jk

OpenStudy (anonymous):

X substitutes values

OpenStudy (hitaro9):

Good, awesome. So X subsitutes numbers. Can it substitute all numbers?

OpenStudy (anonymous):

Yes.....

OpenStudy (hitaro9):

Right, cool. Now, we can plug any number in for X, but not every number will be correct. If I asked you to sub in a couple number in to X > 10 Can you think of a couple examples where that's true, and where that isn't?

OpenStudy (anonymous):

Idk honestly. Im pretty sure it will only substitute certain numbers

OpenStudy (hitaro9):

Okay then, can you expand on that a bit? Why do you think that?

OpenStudy (anonymous):

Because, there are some numbers that won't be able to be substituted. They're to important

OpenStudy (anonymous):

x+15=8 What is the value of X?

pooja195 (pooja195):

Subtract 15 from both sides

OpenStudy (hitaro9):

So X can substitute numbers right? How about we try a few? Guess and check some numbers for me.

OpenStudy (anonymous):

That's my question. How would I answer it? Can you walk me through step by step?

OpenStudy (anonymous):

7

OpenStudy (hitaro9):

Okay, how'd you get 7?

pooja195 (pooja195):

8-15=?

OpenStudy (anonymous):

15-8

pooja195 (pooja195):

no 8-15

OpenStudy (hitaro9):

Okay, that's good, we're starting to see relationships in the equations.

OpenStudy (anonymous):

-7 sorry I did it backwards

pooja195 (pooja195):

its ok thats correct :)

OpenStudy (hitaro9):

With these problems we can always plug in and check to see how we did.

OpenStudy (anonymous):

So that would be my answer?

pooja195 (pooja195):

You did steps wrong

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

What are the steps?

pooja195 (pooja195):

Subtract 12x from both sides. \[\huge~\rm~10x+4−12x=12x−16−12x\]What are you left with ? \[\huge~\rm~−2x+4=−16\]

pooja195 (pooja195):

Now you can solve for x

OpenStudy (hitaro9):

So Durgamata, if I asked you to find 3 numbers that worked for X > 10, and 3 numbers that didn't work, how would you go about that?

OpenStudy (anonymous):

Subtract

pooja195 (pooja195):

Subtract what?

pooja195 (pooja195):

−2x+4=−16 Can you solve for x from here? Try?

OpenStudy (hitaro9):

Durgamata is at a point where they're just starting to understand how variables work. If they're struggling with x+3 =5, I don't think they're going to be able to solve multistep equations, and walking through it probably isn't going to leave them with much that was meaningful

OpenStudy (anonymous):

@pooja195 Im not on that question right now. I had to restart my test and it erased all the answers you helped me with.

pooja195 (pooja195):

Your test?

OpenStudy (anonymous):

Not test. Work

pooja195 (pooja195):

Ok well post the question you are on.

OpenStudy (hitaro9):

Okay, Durgamata, I'm going to draw a comparison here. Do you know what a scale is?

OpenStudy (hitaro9):